# Bohr–Sommerfeld rules for systems

### v1

@@ Abstract.
We present a complete, self-contained formulation of the Bohr–Sommerfeld quantization rule for a semiclassical self-adjoint $2\times 2$ system on the real line, arising from a simple closed curve in phase space. We focus on the case where the principal symbol exhibits eigenvalue crossings within the domain enclosed by the curve—a situation commonly encountered in Dirac-type operators. Building on earlier work on scalar Bohr–Sommerfeld rules and semiclassical treatments of the Harper operator near rational flux quanta, we identify additional contributions to the quantization condition, and derive concise expressions for general self-adjoint $2\times 2$ systems. The resulting formulas give explicit geometric phase corrections and clarify when these phases take quantized values.
@@ 1. Introduction
and we see that there is an eigenvalue crossing if and only if $P$ vanishes somewhere. Let $\mu$ be one of $\lambda_{\pm}$, and $E\in\mathbb{R}$ an energy level. The goal of this paper is to present a complete and self-contained formulation of the Bohr–Sommerfeld rule coming from a simple closed curve $\gamma=\mu^{-1}(E)$ in phase space. We assume $d\mu\neq 0$ and $P\neq 0$ in a neighborhood of $\gamma$, but allow $P$ to vanish inside the domain $D$ enclosed by $\gamma=\partial D$. This behavior is typical for Dirac-type systems. One motivation for our work is to develop a framework for understanding the occurrence of almost flat bands in the spectrum of a Dirac-Harper model for strained moiré lattices, introduced by Timmel and Mele [20] and discussed in Subsection 5.3. In doing so, we build on earlier work on scalar Bohr–Sommerfeld rules by Helffer-Robert [9, 10, 11] together with Helffer-Sjöstrand’s semiclassical treatments of the Harper operator [12, 13]. In [13], Helffer-Sjöstrand studied general systems occurring in their analysis near rational fluxes, but their study does not cover general $H_{0}$ of the form above. We therefore revisit their study and identify additional contributions to the quantization condition and derive concise expressions for general self-adjoint $2\times 2$ systems.
If $P\neq 0$ it is natural to diagonalize $H_{0}$ and reduce to the scalar case. If $P=0$ somewhere inside $D$, we first remove the eigenvalue crossings by slightly perturbing the components of $P$, see Lemma 2.1. We also modify $H_{0}$ away from a simply connected neighborhood of $\gamma$ to ensure $H_{0}$ is bounded, see Lemma 2.2. Both these modifications can be done in a way that doesn’t change the spectrum of $H^{w}(x,hD)$ modulo $\mathcal{O}(h^{\infty})$, see Proposition 2.3. We can then find unitary $U\in S(1)$ such that
@@
contains the Maslov index $\pi$. Here $f_{1}$ is the subprincipal symbol of $\mu^{w}+hD_{11}^{w}$, and the integral $\int_{\gamma}f_{1}\,dt$ contains the geometric phase corrections in the Bohr–Sommerfeld rule. Each subsequent $S_{j}(E)$, $j\geq 2$ can also be computed using an algorithm due to Colin de Verdière [3].
@@
whose off-diagonal entries are the annihilation and creation operators of the harmonic oscillator. This simple model can be seen as a special case of the Jackiw-Reebi model studied in Subsection 5.1. It is known that the operator eigenvalues of (5) coincide with the roots to $e^{iS_{0}(E)/h}=1$, where $S_{0}$ is given by (3). (The corresponding eigenfunctions can be computed explicitly.) The eigenvalues of the principal symbol are $\lambda_{\pm}(x,\xi)=\pm\sqrt{\xi^{2}+x^{2}}$, and the level sets $\lambda_{\pm}=E$ are circles with radius $\lvert E\rvert$. Hence, $S_{0}(E)=\pi E^{2}$ which by (2) gives
@@
This kind of analysis has been used by Helffer and Sjöstrand in their treatment of the Harper operator (see [13, Section 3.6] in particular for operators such as (5)), but, as mentioned, their study does not cover general $H_{0}$ of the form considered here. After diagonalizing the operator $H^{w}(x,hD)$ we find that an additional term appears in the subprincipal symbol which is absent in the Harper setting, see (10). As in [12, 13] we note that the analysis presented here generalizes e.g. to the case when $H_{0}$ is periodic and to many other situations where $\mu^{-1}(E_{0})$ is not just one simple closed curve but instead a countable union of connected components, provided these components are separated by barriers, see Remark 2.4.
Before stating our main results, which include concise expressions for the geometric phase corrections of the Bohr–Sommerfeld rule, together with information about when these phases only take a discrete set of values, i.e., when they become quantized, we briefly note the recent work of Yoshida [21], where a different and inherently non-scalar method was used to obtain the leading-order term in the Bohr–Sommerfeld rule for operators with $p_{2}=\xi$, $p_{3}=V(x)$, and $p_{0}=p_{1}=0.$ The new ingredient in Yoshida’s method is that it applies when $V(x)\in C^{\infty}$ is not necessarily analytic. When $V(x)$ is analytic, the exact WKB method had previously been successfully used to obtain Bohr–Sommerfeld rules for various $2\times 2$ systems, including some non-self-adjoint cases [5, 6, 7, 14, 15]. We also mention that the Bohr–Sommerfeld rules we present below are for energies away from the eigenvalue crossing point—for results on spectrum for energies near these critical points, where the eigenvalues of the principal symbol coalesce, we refer for example to [1, 4, 13] and the references therein. It follows from [12] (see also [16] for a recent treatment of a more general case) that, for scalar operators with a non-degenerate principal symbol near the bottom of the well, the Bohr–Sommerfeld quantization rule remains valid at the bottom of the spectrum. In the case of systems, this applies to operators such as (5), at the eigenvalue crossing, and to their generalizations studied in §3.6 of [13], but it may fail for general $2\times 2$ systems that we study in this work.
@@ 1.1. Statement of results
    $\mu(x,\xi)\in I\text{ for }(x,\xi)\in\mathcal{A}\quad\Longrightarrow\quad\lambda_{-}(x,\xi)<E_{-}.$ (6)
@@
    $\mu(x,\xi)\in I\text{ for }(x,\xi)\in\mathcal{A}\quad\Longrightarrow\quad\lambda_{+}(x,\xi)>E_{+}.$ (7)
@@
$f_{1}=r_{0}\pm\sum_{i=i}^{3}\frac{r_{i}p_{i}}{\lVert P\rVert}+\mu_{1}$
@@
$\mu_{1}=\frac{1}{2i}\big{\langle}\{H_{0}-\mu,e\},e\big{\rangle}+\frac{1}{i}\big{\langle}\{\mu,e\},e\big{\rangle}+\mu\operatorname{Im}\langle e_{x}^{\prime},e_{\xi}^{\prime}\rangle.$ (10)
@@
In [13], the first two terms on the right of (10) are denoted by $\mu_{1}^{\prime}$ and $\mu_{1}^{\prime\prime}$ respectively, while the third term doesn’t appear for the Harper operator, see formula $(6.2.19)$ in [13]. (As shown by Theorem 4.4 the third term vanishes identically under certain symmetry assumptions.) Accordingly, we denote these three terms by $\mu_{1}^{\prime},\mu_{1}^{\prime\prime}$, and $\mu_{1}^{\prime\prime\prime}$. Then $\mu_{1}^{\prime\prime}$ is the Berry connection, which upon integration gives rise to the Berry phase
@@
see Remark 3.4. The terms $\mu_{1}^{\prime}$ and $\mu_{1}^{\prime\prime\prime}$ are related to the Berry curvature scalar density, and upon integration they give rise to the Rammal-Wilkinson phase [^1] While the term Rammal-Wilkinson phase seems to have been reserved for the non-Berry phase correction to the Bohr–Sommerfeld rule for the Harper operator near rational magnetic fluxes [8], we decided to keep this terminology in our setting as well—though the corrections here may include contributions not present in the Harper case.
$\theta_{\mathit{RW}}=\int_{\gamma}(\mu_{1}^{\prime}+\mu_{1}^{\prime\prime\prime})\,dt,$
@@
$H_{0}(x,\xi)=p_{0}(x,\xi)I+\lVert P(x,\xi)\rVert U,\quad U:=\begin{pmatrix}\cos\theta&e^{-i\phi}\sin\theta\\ e^{i\phi}\sin\theta&-\cos\theta\end{pmatrix}$, (12)
where $U$ is unitary and Hermitian.
@@
$\theta_{B}$ $=\pm\int_{\gamma}\frac{1-\cos(\theta)}{2}\{\lambda_{\pm},\phi\}\,dt,\quad\theta_{\mathit{RW}}=\int_{\gamma}(\pm p_{0}+3\lVert P\rVert)\frac{\sin(\theta)}{4}\{\theta,\phi\}\,dt,$
@@
As we can see from Theorem 1.1, unless there are restrictions on the subprincipal symbol $H_{1}$, the term $\int_{\gamma}f_{1}\,dt$ will not be quantized in general. However, in Section 4 we show that the Berry and Rammal-Wilkinson phases are quantized as soon as the components of $P$ are linearly dependent over $\mathbb{R}$, see Theorem 4.4. In particular, this happens when (at least) one of $p_{1},p_{2},p_{3}$ vanishes identically near $\gamma$. The following theorem therefore complements Theorem 1.2, and together they yield a comprehensive description of the general situation.
@@
$2\pi kh=-S_{0}(E)+\sum_{j=1}^{\infty}h^{j}S_{j}(E)+\mathcal{O}(h^{\infty}),$
@@
Let $\theta(\pm H_{0})$ and $\phi(\pm H_{0})$ be spherical coordinates defined as in (11) for $\pm H_{0}$. Then $\theta(-H_{0})=\theta(H_{0})$ while $\phi(-H_{0})=-\phi(H_{0})$. If $\mu=\lambda_{\pm}$ then $\mu_{\mp}$ has a well near the energy level E-E, and $E\in\operatorname{Spec}(-H^{w})\cap[-E_{+},-E_{-}]$ is described by the Bohr–Sommerfeld rule in Corollary 1.4 for $\mu_{\mp}$, with $E$ replaced by E-E and $\gamma$ by $\gamma$ (that is, $\gamma$ is the same curve but oriented in the clockwise direction), while $r_{i},p_{i}$ are replaced by $r_{i},-p_{i}$, and $\theta,\phi$ in the formulas for $\theta_{B},\theta_{\mathit{RW}}$ are replaced by $\theta,-\phi$. Now
$S_{0}(-E)=\int_{\{\mu_{\mp}=-E\};\text{ clockwise}}\xi\,dx=\int_{-\gamma}\xi\,dx=-\int_{\gamma}\xi\,dx,$
@@
$S_{1}(-E)=\pi-\int_{-\gamma}\bigg{(}-r_{0}\mp\sum_{i=1}^{3}\frac{(-r_{i})(-p_{i})}{\lVert P\rVert}\bigg{)}\,dt-(\theta_{B}+\theta_{\mathit{RW}})(\theta,-\phi,\mu_{\mp},-\gamma)$
where we wrote $(\theta_{B}+\theta_{\mathit{RW}})(\theta,-\phi,\mu_{\mp},-\gamma)$ to indicate $\theta_{B}$ and $\theta_{\mathit{RW}}$ should be calculated for eigenvalue $\mu_{\mp}=-p_{0}\mp\lVert P\rVert$ and spherical coordinates $\theta,-\phi$ along $\gamma$. Since $\mu_{\mp}=-\lambda_{\pm}$ we see from Theorems 1.2 and 1.3 that
$\theta_{\bullet}(\theta,-\phi,\mu_{\mp},-\gamma)=\theta_{\bullet}(\theta,\phi,\lambda_{\pm},\gamma),\quad\bullet=B,\mathit{RW},$
which gives
$S_{1}(-E)=\pi-\int_{\gamma}\bigg{(}r_{0}\pm\sum_{i=1}^{3}\frac{r_{i}p_{i}}{\lVert P\rVert}\bigg{)}\,dt-\theta_{B}-\theta_{\mathit{RW}}.$
with $\theta_{B}$ and $\theta_{\mathit{RW}}$ as in Theorems 1.2 and 1.3. Hence, if we define $S_{0}(E)$ and $S_{1}(E)$ by the same formal expressions as in the well case, then $S_{0}(E)=-S_{0}(-E)$ and $S_{1}(E)=S_{1}(-E)$. Using the algorithm of Colin de Verdière [3] applied to $H^{w}$ it follows from Remark 1 and Theorem 1 in the mentioned paper that $S_{j}(-E)=S_{j}(E)$ for $j\geq 2$ as well if we define $S_{j}$ by the same formal expressions. ∎
@@ 2. Microlocal preparation
We give the proof in the case when $\mu=\lambda_{+}$, and use (6). The case $\mu=\lambda_{-}$ is similar, see (7). Since $\mu(A_{+})=E_{+}$ we can take a cutoff function $\psi\in C_{c}^{\infty}$ with $\psi=1$ in a neighborhood of $W$ where $\lambda_{+}>\lambda_{-}$ (possibly after using Lemma 2.1 to remove zeros of $P$ inside $D$), such that if $w\in\operatorname{supp}\nabla\psi$ then $\lambda_{+}(x,\xi)\geq E_{+}+\varepsilon$ and $\lambda_{-}(x,\xi)\leq E_{-}-\varepsilon$ for some $\varepsilon>0$. Introduce modified spherical coordinates $\widetilde{\theta}=\operatorname{arccos}(\psi p_{3}/\lVert P\rVert)$ and $\widetilde{\phi}=\operatorname{sgn}(p_{2})\operatorname{arcccos}(\psi p_{1}/\sqrt{p_{1}^{2}+p_{2}^{2}})$, and define $\widetilde{U}$ as $U$ in (12) but with $\theta,\phi$ replaced by $\widetilde{\theta},\widetilde{\phi}$. Set
$\widetilde{H}_{0}=\psi p_{0}+(1-\psi)\frac{E_{+}+E_{-}}{2}+\bigg{(}\psi\lVert P\rVert+(1-\psi)\bigg{(}\frac{E_{+}-E_{-}}{2}+\varepsilon\bigg{)}\bigg{)}\widetilde{U}$.
@@
for some $k\times k$ matrix-valued $A_{11}$ and $(n-k)\times(n-k)$ matrix-valued $A_{22}$. If $A_{11}(x,\xi)$ and $A_{22}(x,\xi)$ have disjoint sets of eigenvalues, uniformly for $(x,\xi)\in T^{*}\mathbb{R}^{d}$, then there is a unitary $U\sim\sum h^{j}U_{j}$ with $U_{j}\in S(1)$ such that
$(U^{w})^{*}M^{w}U^{w}=\begin{pmatrix}A_{11}^{w}&\\ &A_{22}^{w}\end{pmatrix}+hD^{w}+\mathcal{\mathcal{missing}}O_{L^{2}\to L^{2}}(h^{\infty})$
@@
This is just a reformulation of [13, Proposition 3.1.1] and [13, Corollary 3.1.2] for the Weyl quantization, and follows by using Taylor’s trick [18] adapted to semiclassical operators with an additional argument to make $U$ unitary (see [13, Corollary 3.1.2]). Taylor’s approach ensures that there are no terms $U_{1}^{*}M_{0}U_{0}+U_{0}^{*}M_{0}U_{1}$ in the principal symbol of $D$, so the last formula follows directly from the Weyl calculus. ∎
We apply the theorem to $H$ with $A_{11}=\mu$, where $\mu\in\{\lambda_{+},\lambda_{-}\}$ is the eigenvalue satisfying (8), and $A_{22}$ is the other one. We define $U_{0}$ using a pair of orthonormal eigenvectors, and let $e$ be the eigenvector corresponding to $\mu$ so that the first column of $U_{0}$ is $e$. By the theorem, we then have
@@
$f_{1}=\Big{(}U_{0}^{*}H_{1}U_{0}+\frac{1}{2i}\Big{(}U_{0}^{*}\{H_{0},U_{0}\}+\{U_{0}^{*},H_{0}U_{0}\}\Big{)}\Big{)}_{11}=(U_{0}^{*}H_{1}U_{0})_{11}+\mu_{1}.$ (13)
Here $\mu_{1}$ is the top left entry of the subprincipal symbol of $(U_{0}^{w})^{*}H_{0}^{w}U_{0}^{w}$. Since the top row of $U_{0}^{*}$ is $\bar{e}$, and the first column of a matrix product $AB$ is $A$ times the first column of $B$, we get
@@
$\mu_{1}$ $=\frac{1}{2i}\Big{[}\bar{e}\cdot((H_{0})_{\xi}^{\prime}e_{x}^{\prime}-(H_{0})_{x}^{\prime}e_{\xi}^{\prime})+\{\bar{e},\mu e\}\Big{]}$
$=\frac{1}{2i}\Big{[}\big{\langle}\{H_{0},e\},e\big{\rangle}+\langle\partial_{x}(\mu e),\partial_{\xi}e\rangle-\langle\partial_{\xi}(\mu e),\partial_{x}e\rangle\Big{]}.$
Expanding the derivatives of $\mu e$, and using the fact that $\langle e,e\rangle=1$ implies upon differentiating that
@@
we get
$\mu_{1}=\frac{1}{2i}\Big{[}\big{\langle}\{H_{0}+\mu,e\},e\big{\rangle}+\mu\big{(}\langle e_{x}^{\prime},e_{\xi}^{\prime}\rangle-\langle e_{\xi}^{\prime},e_{x}^{\prime}\rangle\big{)}\Big{]}.$
We rewrite this as
$\mu_{1}=\frac{1}{2i}\big{\langle}\{H_{0}-\mu,e\},e\big{\rangle}+\frac{1}{i}\big{\langle}\{\mu,e\},e\big{\rangle}+\mu\operatorname{Im}\langle e_{x}^{\prime},e_{\xi}^{\prime}\rangle,$
which is formula (10) from the introduction. Accordingly, we denote the three terms on the right by $\mu_{1}^{\prime},\mu_{1}^{\prime\prime}$, and $\mu_{1}^{\prime\prime\prime}$. By formula $(6.2.30)$ in [13] it follows that the curve integrals of $\mu_{1}^{\prime}$ and $\mu_{1}^{\prime\prime}$ are intrinsically defined in terms of the spectral projection $\Pi$ associated to $\mu$ and don’t depend on the choice of eigenvector. The same is true for $\mu_{1}^{\prime\prime\prime}$. Indeed, let $\Pi v=\langle v,e\rangle e$. A straightforward calculation then shows that
[…]
which proves the claim. (We discuss this in greater detail in Lemmas 3.5 and 3.7 below.) Since $d\Pi=\Pi_{x}^{\prime}dx+\Pi_{\xi}^{\prime}d\xi$, we have that
$d\Pi\wedge d\Pi=[\Pi_{x}^{\prime},\Pi_{\xi}^{\prime}]\,dx\wedge d\xi$
and thus
[…]
The quantity
[…]
is the Berry curvature.

### v2

@@ Abstract.
We present a complete, self-contained formulation of the Bohr–Sommerfeld quantization rule for a semiclassical self-adjoint $2\times 2$ system on the real line, arising from a simple closed curve in phase space. We focus on the case where the principal symbol exhibits eigenvalue crossings within the domain enclosed by the curve – a situation commonly encountered in Dirac-type operators. Building on earlier work on scalar Bohr–Sommerfeld rules and semiclassical treatments of the Harper operator near rational flux quanta, we derive concise expressions for general self-adjoint $2\times 2$ systems. The resulting formulas give explicit geometric phase corrections and clarify when these phases take quantized values.
@@ 1. Introduction
and we see that there is an eigenvalue crossing if and only if $P$ vanishes somewhere.
Let $\mu$ be one of $\lambda_{\pm}$, and $E\in\mathbb{R}$ an energy level. The goal of this paper is to present a complete and self-contained formulation of the Bohr–Sommerfeld rule coming from a simple closed curve $\gamma=\mu^{-1}(E)$ in phase space. We assume $d\mu\neq 0$ and $P\neq 0$ in a neighborhood of $\gamma$, but allow $P$ to vanish inside the domain $D$ enclosed by $\gamma=\partial D$. This behavior is typical for Dirac-type systems. One motivation for our work is to develop a framework for understanding the occurrence of almost flat bands in the spectrum of a Dirac-Harper model for strained moiré lattices, introduced by Timmel and Mele [20] and discussed in Subsection 5.3. In doing so, we build on earlier work on scalar Bohr–Sommerfeld rules by Helffer-Robert [9, 10, 11] together with Helffer-Sjöstrand’s semiclassical treatments of the Harper operator [12, 13] near rational fluxes, which we revisit to derive concise expressions for general self-adjoint $2\times 2$ systems.
If $P\neq 0$ it is natural to diagonalize $H_{0}$ and reduce to the scalar case. If $P=0$ somewhere inside $D$, we first remove the eigenvalue crossings by slightly perturbing the components of $P$, see Lemma 2.1. We also modify $H_{0}$ away from a simply connected neighborhood of $\gamma$ to ensure $H_{0}$ is bounded, see Lemma 2.2. Both these modifications can be done in a way that doesn’t change the spectrum of $H^{w}(x,hD)$ modulo $\mathcal{O}(h^{\infty})$, see Proposition 2.3. We can then find $U\in S(1)$ such that $U^{w}$ is unitary and
@@
contains the Maslov index in the $\pi$ term. Here $f_{1}$ is the subprincipal symbol of $\mu^{w}+hD_{11}^{w}$, and the integral $\int_{\gamma}f_{1}\,dt$ contains the geometric phase corrections in the Bohr–Sommerfeld rule. Each subsequent $S_{j}(E)$, $j\geq 2$ can also be computed using an algorithm due to Colin de Verdière [3].
@@
whose off-diagonal entries are the annihilation and creation operators of the harmonic oscillator. This simple model can be seen as a special case of the Jackiw-Rebbi model studied in Subsection 5.1. It is known that the operator eigenvalues of (5) coincide with the roots to $e^{iS_{0}(E)/h}=1$, where $S_{0}$ is given by (3). (The corresponding eigenfunctions can be computed explicitly.) The eigenvalues of the principal symbol are $\lambda_{\pm}(x,\xi)=\pm\sqrt{\xi^{2}+x^{2}}$, and the level sets $\lambda_{\pm}=E$ are circles with radius $\lvert E\rvert$. Hence, $S_{0}(E)=\pi E^{2}$ which by (2) gives
@@
This kind of analysis has been used by Helffer and Sjöstrand in their treatment of the Harper operator (see [13, Section 3.6] in particular for operators such as (5)). As in [12, 13] we note that the analysis presented here generalizes e.g. to the case when $H_{0}$ is periodic and to many other situations where $\mu^{-1}(E_{0})$ is not just one simple closed curve but instead a countable union of connected components, provided these components are separated by barriers, see Remark 2.4.
Before stating our main results, which include concise expressions for the geometric phase corrections of the Bohr–Sommerfeld rule, together with information about when these phases only take a discrete set of values, i.e., when they become quantized, we briefly note the recent work of Yoshida [21], where a different and inherently non-scalar method was used to obtain the leading-order term in the Bohr–Sommerfeld rule for operators with $p_{2}=\xi$, $p_{3}=V(x)$, and $p_{0}=p_{1}=0.$ The new ingredient in Yoshida’s method is that it applies when $V(x)\in C^{\infty}$ is not necessarily analytic. When $V(x)$ is analytic, the exact WKB method had previously been successfully used to obtain Bohr–Sommerfeld rules for various $2\times 2$ systems, including some non-self-adjoint cases [5, 6, 7, 14, 15]. We also mention that the Bohr–Sommerfeld rules we present below are for energies away from the eigenvalue crossing point – for results on spectrum for energies near these critical points, where the eigenvalues of the principal symbol coalesce, we refer for example to [1, 4, 13] and the references therein. It follows from [12] (see also [16] for a recent treatment of a more general case) that, for scalar operators with a non-degenerate principal symbol near the bottom of the well, the Bohr–Sommerfeld quantization rule remains valid at the bottom of the spectrum. In the case of systems, this applies to operators such as (5), at the eigenvalue crossing, and to their generalizations studied in §3.6 of [13], but it may fail for general $2\times 2$ systems that we study in this work.
@@ 1.1. Statement of results
    $\lambda_{-}(x,\xi)<E_{-}\quad\text{ for all }(x,\xi)\in\mathcal{A}.$ (6)
@@
    $\lambda_{+}(x,\xi)>E_{+}\quad\text{ for all }(x,\xi)\in\mathcal{A}.$ (7)
@@
$f_{1}=r_{0}\pm\sum_{i=1}^{3}\frac{r_{i}p_{i}}{\lVert P\rVert}+\mu_{1}$
@@
$\mu_{1}=\frac{1}{2i}\big\langle\{H_{0}-\mu,e\},e\big\rangle+\frac{1}{i}\big\langle\{\mu,e\},e\big\rangle.$ (10)
@@
In [13], the two terms on the right of (10) are denoted by $\mu_{1}^{\prime}$ and $\mu_{1}^{\prime\prime}$ respectively, see formula $(6.2.19)$ in [13]. Accordingly, we denote them by $\mu_{1}^{\prime}$ and $\mu_{1}^{\prime\prime}$ as well. Then $\mu_{1}^{\prime\prime}$ is the Berry connection, which upon integration gives rise to the Berry phase
@@
see Remark 3.3. The term $\mu_{1}^{\prime}$ is related to the Berry curvature scalar density, and upon integration gives rise to the Rammal–Wilkinson phase
$\theta_{\mathit{RW}}=\int_{\gamma}\mu_{1}^{\prime}\,dt,$
@@
$H_{0}=p_{0}I+\lVert P\rVert R,\quad R:=\begin{pmatrix}\cos\theta&e^{-i\phi}\sin\theta\\ e^{i\phi}\sin\theta&-\cos\theta\end{pmatrix}$, (12)
where $R$ is unitary and Hermitian.
@@
$\theta_{B}$ $=\pm\int_{\gamma}\frac{1-\cos(\theta)}{2}\{\lambda_{\pm},\phi\}\,dt,\quad\theta_{\mathit{RW}}=\int_{\gamma}\lVert P\rVert\frac{\sin(\theta)}{2}\{\theta,\phi\}\,dt,$
@@
As we can see from Theorem 1.1, unless there are restrictions on the subprincipal symbol $H_{1}$, the term $\int_{\gamma}f_{1}\,dt$ will not be quantized in general. However, in Section 4 we show that the Berry and Rammal–Wilkinson phases are quantized as soon as the components of $P$ are linearly dependent over $\mathbb{R}$, see Theorem 4.4. In particular, this happens when (at least) one of $p_{1},p_{2},p_{3}$ vanishes identically near $\gamma$. The following theorem therefore complements Theorem 1.2, and together they yield a comprehensive description of the general situation.
@@
$2\pi kh=-S_{0}(E)+\sum_{j=1}^{\infty}S_{j}(E)h^{j}+\mathcal{O}(h^{\infty}),$
@@
If $\mu=\lambda_{\pm}$ then $\mu_{\mp}$ has a well near the energy level E-E, and $E\in\operatorname{Spec}(-H^{w})\cap[-E_{+},-E_{-}]$ is described by the Bohr–Sommerfeld rule in Theorem 1.1 for $\mu_{\mp}$, with $E$ replaced by E-E, $\gamma$ by $\gamma$ (that is, $\gamma$ is the same curve but oriented in the clockwise direction), and $f_{1}$ by $f_{1}$. Now
$S_{0}^{\mathrm{well}}(-E)=\int_{\{\mu_{\mp}=-E\};\text{ clockwise}}\xi\,dx=\int_{-\gamma}\xi\,dx=-\int_{\gamma}\xi\,dx,$
@@
$S_{1}^{\mathrm{well}}(-E)=\pi-\int_{-\gamma}(-f_{1})\,dt=\pi-\int_{\gamma}f_{1}\,dt.$
Hence, if we define $S_{0}^{\mathrm{barrier}}(E)$ and $S_{1}^{\mathrm{barrier}}(E)$ by the same formal expressions as in the well case, so that $S_{0}^{\mathrm{barrier}}(E)=\int_{\gamma}\xi\,dx$ and $S_{1}^{\mathrm{barrier}}(E)=\pi-\int_{\gamma}f_{1}\,dt$, then
$2\pi kh=-S_{0}^{\mathrm{barrier}}(E)+hS_{1}^{\mathrm{barrier}}(E)+\sum_{j=2}^{\infty}S_{j}^{\mathrm{well}}(-E)h^{j}.$
Using the algorithm of Colin de Verdière [3] applied to $H^{w}$ it follows from Remark 1 and Theorem 1 in the mentioned paper that $S_{j}^{\mathrm{well}}(-E)=S_{j}^{\mathrm{barrier}}(E)$ for $j\geq 2$ too if we define $S_{j}^{\mathrm{barrier}}$ by the same formal expressions as in the case of a well. ∎
@@ 2. Microlocal preparation
We give the proof in the case when $\mu=\lambda_{+}$, and use (6). The case $\mu=\lambda_{-}$ is similar, see (7). Since $\mu(A_{+})=E_{+}$ we can take a cutoff function $\psi\in C_{c}^{\infty}$ with $\psi=1$ in a neighborhood of $W$ where $\lambda_{+}>\lambda_{-}$ (possibly after using Lemma 2.1 to remove zeros of $P$ inside $D$), such that if $w\in\operatorname{supp}\nabla\psi$ then $\lambda_{+}(x,\xi)\geq E_{+}+\varepsilon$ and $\lambda_{-}(x,\xi)\leq E_{-}-\varepsilon$ for some $\varepsilon>0$. Introduce modified spherical coordinates $\widetilde{\theta}=\operatorname{arccos}(\psi p_{3}/\lVert P\rVert)$ and $\widetilde{\phi}=\operatorname{sgn}(p_{2})\operatorname{arcccos}(\psi p_{1}/\sqrt{p_{1}^{2}+p_{2}^{2}})$, and define $\widetilde{R}$ as $R$ in (12) but with $\theta,\phi$ replaced by $\widetilde{\theta},\widetilde{\phi}$. Set
$\widetilde{H}_{0}=\psi p_{0}+(1-\psi)\frac{E_{+}+E_{-}}{2}+\bigg(\psi\lVert P\rVert+(1-\psi)\bigg(\frac{E_{+}-E_{-}}{2}+\varepsilon\bigg)\bigg)\widetilde{R}$.
@@
for some $k\times k$ matrix-valued $A_{11}$ and $(n-k)\times(n-k)$ matrix-valued $A_{22}$. If $A_{11}(x,\xi)$ and $A_{22}(x,\xi)$ have disjoint sets of eigenvalues, uniformly for $(x,\xi)\in T^{*}\mathbb{R}^{d}$, then there is a unitary $U^{w}\sim\sum h^{j}U_{j}^{w}$ with $U_{j}\in S(1)$ such that
$(U^{w})^{*}M^{w}U^{w}=\begin{pmatrix}A_{11}^{w}&\\ &A_{22}^{w}\end{pmatrix}+hD^{w}+\mathcal{\mathcal{}}O_{L^{2}\to L^{2}}(h^{\infty})$
@@
This is just a reformulation of [13, Proposition 3.1.1] and [13, Corollary 3.1.2] for the Weyl quantization, and follows by using Taylor’s trick [18] adapted to semiclassical operators with an additional argument to make $U^{w}$ unitary. The last formula follows directly from the Weyl calculus. ∎
We apply the theorem to $H$ with $A_{11}=\mu$, where $\mu\in\{\lambda_{+},\lambda_{-}\}$ is the eigenvalue satisfying (8), and $A_{22}$ is the other one. Since $U^{w}$ is unitary, and $(U^{w})^{*}=(U^{*})^{w}$, the subprincipal symbol of $I=(U^{*})^{w}U^{w}$ has to vanish. In view of the Weyl calculus, this means that
$U_{1}^{*}U_{0}+U_{0}^{*}U_{1}+\frac{1}{2i}\{U_{0}^{*},U_{0}\}=0.$
Hence,
$(U_{1}^{*}H_{0}U_{0}+U_{0}^{*}H_{0}U_{1})_{11}=\mu(U_{1}^{*}U_{0}+U_{0}^{*}U_{1})_{11}=-\frac{\mu}{2i}\{U_{0}^{*},U_{0}\}_{11}.$
We define $U_{0}$ using a pair of orthonormal eigenvectors, and let $e$ be the eigenvector corresponding to $\mu$ so that the first column of $U_{0}$ is $e$. By the theorem, we then have
@@
$f_{1}$ $=\Big(U_{0}^{*}H_{1}U_{0}+\frac{1}{2i}\Big(U_{0}^{*}\{H_{0},U_{0}\}+\{U_{0}^{*},H_{0}U_{0}\}-\mu\{U_{0}^{*},U_{0}\}\Big)\Big)_{11}$ (13)
$=(U_{0}^{*}H_{1}U_{0})_{11}+\mu_{1}.$
Here $\mu_{1}$ is the top left entry of the subprincipal symbol of $(U^{w})^{*}H_{0}^{w}U^{w}$. Since the top row of $U_{0}^{*}$ is $\bar{e}$, and the first column of a matrix product $AB$ is $A$ times the first column of $B$, we get
@@
$\mu_{1}$ $=\frac{1}{2i}\Big[\bar{e}\cdot((H_{0})_{\xi}^{\prime}e_{x}^{\prime}-(H_{0})_{x}^{\prime}e_{\xi}^{\prime})+\{\bar{e},\mu e\}-\mu\{\bar{e},e\}\Big]$
$=\frac{1}{2i}\Big[\big\langle\{H_{0},e\},e\big\rangle+\mu_{x}^{\prime}\langle e,e_{\xi}^{\prime}\rangle-\mu_{\xi}^{\prime}\langle e,e_{x}^{\prime}\rangle\Big].$
Differentiating $\langle e,e\rangle=1$ gives
@@
so $\mu_{1}=\frac{1}{2i}\big\langle\{H_{0}+\mu,e\},e\big\rangle$. We rewrite this as
$\mu_{1}=\frac{1}{2i}\big\langle\{H_{0}-\mu,e\},e\big\rangle+\frac{1}{i}\big\langle\{\mu,e\},e\big\rangle,$
which is formula (10) from the introduction. Accordingly, we denote the two terms on the right by $\mu_{1}^{\prime}$ and $\mu_{1}^{\prime\prime}$. This also recovers formula $(6.2.19)$ in [13] using the Weyl calculus instead of the calculus for the standard quantization. By formula $(6.2.30)$ in [13] it follows that the curve integrals of $\mu_{1}^{\prime}$ and $\mu_{1}^{\prime\prime}$ are intrinsically defined in terms of the spectral projection $\Pi$ associated to $\mu$ and don’t depend on the choice of eigenvector. (We discuss this in greater detail in Lemmas 3.4 and 3.6 below.)